Trichloroethylene (TCE) is a common solvent used to degrease machine parts. calculate the empirical formula for TCE if the percent composition is 18.25% C, 0.77% H, and 80.99% Cl.
Answer #1

we have mass of each elements as:
C: 18.25 g
H: 0.77 g
Cl: 80.99 g


Divide by molar mass to get number of moles of each:
C: 18.25/12.01 = 1.5196
H: 0.77/1.008 = 0.7639
Cl: 80.99/35.45 = 2.2846


Divide by smallest to get simplest whole number ratio:
C: 1.5196/0.7639 = 2
H: 0.7639/0.7639 = 1
Cl: 2.2846/0.7639 = 3


So empirical formula is:C2HCl3
Answer: C2HCl3